### v2

# Content Conditional Debiasing for Fair Text Embedding

Abstract
Mitigating biases in machine learning models has gained increasing attention in Natural Language Processing (NLP). Yet, only a few studies focus on fair text embeddings, which are crucial yet challenging for real-world applications. In this paper, we propose a novel method for learning fair text embeddings. We achieve fairness while maintaining utility trade-off by ensuring conditional independence between sensitive attributes and text embeddings conditioned on the content. Specifically, we enforce that embeddings of texts with different sensitive attributes but identical content maintain the same distance toward the embedding of their corresponding neutral text. Furthermore, we address the issue of lacking proper training data by using Large Language Models (LLMs) to augment texts into different sensitive groups. Our extensive evaluations demonstrate that our approach effectively improves fairness while preserving the utility of embeddings, representing a pioneering effort in achieving conditional independence for fair text embeddings.
Content Conditional Debiasing for Fair Text Embedding
[…]
1 Introduction
Embedding text into dense representations is popular in modern Natural Language Processing (NLP), enables applications like sentiment analysis Dang et al. (2020), recommendation systems Zhang et al. (2016), and search engines Palangi et al. (2016). However, as these embeddings become widespread, their inherent biases could affect various applications Packer et al. (2018); Baeza-Yates (2018); Zerveas et al. (2022); Rabelo et al. (2022), potentially harming marginalized groups by perpetuating gender, race, and other biases. Unfortunately, various forms of biases, including gender, racial, and religious biases, have been identified in text embeddings generated by pre-trained language models (PLMs), as reported in several studies Bolukbasi et al. (2016); Nissim et al. (2020); Liang et al. (2020); May et al. (2019). Consequently, attaining fairness in text embedding models is crucial.
Figure 1: Pipleline of our method with gender as the sensitive attributes. (a) Graphical demonstration of the fairness issue. (b) The debiasing procedure achieves a content-conditioned equal distance to improve the fairness. (c) Overview of the data augmentation strategy, including the prompt template used to replace sensitive words with their equivalents from all sensitive groups. (d) Prompt search module: Augmented texts are sent to the demographic polarity checking block. Incorrectly augmented samples are then manually labeled and added to the prompts.
Recent debiasing techniques Liang et al. (2020); Kaneko and Bollegala (2021) for text embeddings use post-training to address biases, avoiding the inefficiency of retraining sentence encoders for each new bias. When removing bias, projection-based methods Liang et al. (2020); Kaneko and Bollegala (2021) reduce an embedding’s projection onto each bias subspace. The distance-based method Yang et al. (2023) constructs embeddings for sensitive groups and equalizes distances to text embeddings across these groups. Nevertheless, these methods persist in pursuing independence between sensitive attributes and text embeddings completely remove the sensitive information. As a result, these approaches do not effectively find the sweet spot between fairness and utility trade-off Zhao and Gordon (2022); Deng et al. (2023); Zliobaite (2015).
Recent studies Mary et al. (2019); Deng et al. (2023); Pogodin et al. (2022) suggest that using datasets labeled with sensitive information to achieve conditional independence — specifically, conditioning on the content class to preserve semantic information within the text — provides a more effective approach to achieving fairness while preserving utility. Yet, the scarcity of text datasets with sensitive labels Gallegos et al. (2023) limits the practical application of these findings. Although Counterfactual Data Augmentation (CDA) Zhao et al. (2018) and LLMs Schick and Schütze (2021); Shao et al. (2023) provide ways for generating data, they face challenges with grammatically incorrect artificial instances Dinan et al. (2019) or inherent systematic biases Yu et al. (2023), respectively.
In this paper, we propose a new method to improve text embedding fairness via Content Conditional sensitive attributes Debiasing (dubbed CCD). Specifically, we first use LLMs to augment a dataset into texts across various sensitive groups linked to the same content by substituting sensitive words with equivalents from different groups (i.e., "son", "daughter" and "child" for gender bias). Then, we achieve conditional independence between sensitive attributes and text embeddings by ensuring that texts with varied sensitive attributes but identical content have embeddings equidistant from their neutral counterparts. Extensive evaluations on debiasing benchmarks and downstream tasks demonstrate our method’s effectiveness in promoting fairness while preserving utility.
2 Method
2.1 Problem Setting
This section outlines the problem of fairness in text embeddings. We define key variables: $S$ for input text, $C$ for the content, and $f$ for the text embedding model that generates text embedding $C^{\prime}=f(S)$. We also denote sensitive attributes as $A=[a_{1},...,a_{|A|}]$ (e.g., gender, race, age), words related to attribute $a_{i}$ as $X^{a_{i}}$, and neutral words as $X^{n}$. A text is neutral ($S^{n}$) if it lacks sensitive words, and is associated with attribute $a_{i}$ ($S^{a_{i}}$) based on its sensitive polarity Wang et al. (2023). We list all notations in Appendix Table 3.
Fairness Issue: Fig. 1 (a) shows there exists an association between attributes $A$ and content variable $C$, if model $f$ superficially considers $A$ as a proxy for $C$, leading to issues of fairness.
2.2 Content Conditional Fairness Loss
We propose to achieve conditional independence between sensitive attributes and content $A\perp C^{\prime}~{}|~{}C$. The conditional independence allows prediction $C^{\prime}$ to depend on $A$ but only through the content variable $C$, prohibiting abusing $A$ as a proxy for $C$ thus mitigating the fairness issue while preserving the utility. Echoing the methodologies in previous research Hinton and Roweis (2002); Yang et al. (2023), we can define the probability $P(i|j)$ as the likelihood of variable $i$ selecting $j$ as its neighbor in the embedding space. Then conditional independence $A\perp C^{\prime}~{}|~{}C$ can be represented as a content conditioned equal distance on the embedding space, as shown in Fig. 1 (b) and definition 2.1 (detailed proof in Appendix B.5) :
Definition 2.1.
(Content Conditional Equal Distance Fairness.) Let $S_{C}^{n}$ be a neutral text with the content $C$, assume $S_{C}^{A}=[S_{C}^{a_{1}},S_{C}^{a_{2}},...,S_{C}^{a_{|A|}}]$ being a set of texts from all sensitive groups with the same content $C$. Then embedding model $f$ is content conditioned equal distance fair with respect to attribute $A$, for any $a_{i},a_{j}\in A$:
$||f(S_{C}^{a_{i}})-f(S_{C}^{n})||=||f(S_{C}^{a_{j}})$ $-f(S_{C}^{n})||.$ (1)
To align with this definition of fairness, we design a loss function $L_{bias}$ that aims to mitigate biases while preserving the representation ability of PLMs. For a sample pair $[S_{C}^{a_{1}},...,S_{C}^{a_{|A|}},S_{C}^{n}]$ :
$L_{bias}=\sum_{i\in[A]}\sum_{j\not=i}$ $|dist(f(S_{C}^{a_{i}}),f(S_{C}^{n}))$
$-dist(f(S_{C}^{a_{j}}),f(S_{C}^{n}))|,$ (2)
where $dist(A,B)=\exp\left(-\frac{\lVert A-B\rVert^{2}}{2\rho^{2}}\right)$ measures the distance on the manifold Yang et al. (2023); Hinton and Roweis (2002) (details in Appendix B.5), and $\rho$ is selected as the variance of the distance over the training dataset for normalization. To further preserve the valuable information encoded in the model for downstream tasks, we design $L_{rep}$ to enforce similarity between the text embeddings processed by the fine-tuned model $f$ and those processed by the original model $f^{ori}$:
$L_{rep}=||f(S^{n})-f^{ori}(S^{n})||.$ (3)
Thus, the overall training objective is:
$L_{all}=L_{bias}+\beta*L_{rep}.$ (4)
2.3 Content Conditional Data Augmentation
To meet our goal in definition 2.1, we augment a dataset into texts across various sensitive attributes $A$ linked to the same content variable $C$.
Augment Text into Different Sensitive Groups: We facilitate text augmentation by using a pretrained LLM to replace sensitive words with equivalents from different sensitive groups. As shown in Fig. 1 (a), for sensitive words $X^{A}=[X^{a_{i}},...,X^{a_{j}}],a_{i},a_{j}\in A$ in text $S$, the LLM $h$ substitutes $X^{A}$ with words from different sensitive groups $h(X^{A})=[[X_{a_{m}}^{a_{i}},...,X_{a_{m}}^{a_{j}}]]_{a_{m}\in A}$ respectively, creating augmented texts $S^{A}$ from all sensitive groups. We also generate a neutral version $S^{n}$ using neutral replacements, resulting in a dataset of texts from different groups $\mathcal{S}^{A}$ and a neutral text $S^{n}$. Examples are provided in Appendix Table 4.
Rule Guided Prompt Searching: Considering the impact of few-shot prompt tuning on LLM effectiveness, choosing the right prompts is essential for successful text augmentation. As shown in Fig. 1 (d), we first find inaccurately augmented texts whose augmented text’s sensitive group mismatches its sensitive polarity Wang et al. (2023) computed by:
$g(S)=\arg\max_{a_{i}\in A}occ(S,V^{a_{i}}),$ (5)
where $occ$ counts occurrences of words from the sensitive words lists $V^{a_{i}},a_{i}\in A$ in the sentence $S$, we assign polarity based on the most frequently mentioned group in $A$. A text $S^{a_{i}}$ is considered wrongly augmented if $g(S^{a_{i}})\not=i$. We then manually augment such texts to create example prompts for better alignment and repeat the process. This process is iterated to select $K$ (ours 10) samples as few-shot prompts.
3 Experiments
In this paper, we take gender bias as an example due to its broad impact on society.
Datasets We generate training data with the News-commentary-v15 corpus Tiedemann (2012). For gender bias evaluation, we follow Yang et al. (2023) to use SEAT May et al. (2019), CrowS-Pairs Nangia et al. (2020) and StereoSet-Intrasentence data Nadeem et al. (2020). We additionally assess fairness on longer texts via the Bias-IR dataset Krieg et al. (2023). To evaluate whether the biased models’ representation ability is maintained, we follow Kaneko and Bollegala (2021); Yang et al. (2023) to select four small-scale subsequent tasks from the GLEU benchmark: Stanford Sentiment Treebank (SST-2 Socher et al. (2013)), Microsoft Research Paraphrase Corpus (MRPC Dolan and Brockett (2005)), Recognizing Textual Entailment (RTE Bentivogli et al. (2009)) and Winograd Schema Challenge (WNLI Levesque et al. (2012)).
Baseline Methods To assess debiasing performance, we compare our algorithm with finetuning-based methods DPCE Kaneko and Bollegala (2021) and ADEPT-F Yang et al. (2023). To assess the effectiveness of our data augmentation strategy, we compare our approach with CDA Zhao et al. (2018). For the use of LLMs in data augmentation, we use ChatGPT and Gemini Team et al. (2023).
Hyperparameters We follow Yang et al. (2023) to conduct experiments on the BERT-LARGE-UNCASED pre-trained model. We use Adam to optimize the objective function. During the debiasing process, our learning rate is 5e-5, batch size is 32, and $\beta$ is 1. Results for DPCE and ADEPT-F are with originally reported hyperparameters Kaneko and Bollegala (2021); Yang et al. (2023). Experiments are conducted on a NVIDIA A100 GPU.
3.1 Results and Analysis
Table 1: Evaluation results on debiasing performance. We test the debiased models on SEAT, CrowS-Pairs, and filtered StereoSet-Intrasentence, with the best result in bold.
Datasets Method SEAT $(0.00)$ the best StereoSet:gender StereoSet:all CrowS-Pairs
[…]
BERT $0.37$ $0.20$ $0.42$ $0.22$ $-0.26$ $0.71$ $0.36$ $86.34$ $59.66$ $69.66$ $84.16$ $58.24$ $70.29$ $55.73$
DPCE $-0.21$ $0.27$ $0.44$ $0.07$ $0.25$ $0.21$ 0.24 $81.19$ $56.72$ $65.41$ $64.06$ 52.96 $60.26$ $52.29$
ADEPT-F $0.83$ $-0.14$ $0.63$ $1.24$ $0.43$ $1.28$ $0.76$ $86.45$ $61.70$ $66.21$ $85.09$ $57.52$ $72.26$ $51.91$
CCD-CDA $0.16$ $0.03$ $0.43$ $0.38$ $0.47$ $0.22$ $0.29$ $80.34$ $\mathbf{53.53}$ $74.69$ $79.10$ $\underline{53.46}$ $73.62$ $46.95$
CCD-GPT $0.47$ $-0.06$ $-0.33$ $0.55$ $-0.17$ $0.48$ $0.34$ $83.76$ $55.40$ $\underline{74.71}$ $82.70$ $54.80$ $\mathbf{74.75}$ $\mathbf{49.62}$
CCD-Gemini $0.47$ $-0.00$ $-0.02$ $-0.72$ $-0.30$ $0.07$ $\underline{0.26}$ $82.91$ $\underline{54.93}$ $\mathbf{74.72}$ $82.97$ $55.00$ $\underline{74.67}$ $\underline{48.85}$
Table 2: Evaluation results on the GLUE dataset and the Bias-IR dataset, we additionally calculate the average deviation to 0.5 for Bias-IR as AVG-DEV. The bold and underline represent the best and second-best respectively.
Datasets Method GLUE $\uparrow$ Bias-IR (Male Ratio, 0.50 the best)
[…]
BERT $92.9$ $84.6$ $\underline{72.5}$ $38.0$ $72.0$ $0.71$ $0.50$ $0.75$ $0.46$ $0.75$ $0.68$ $0.61$ $0.64$ $0.16$
DPCE $92.8$ $69.6$ $53.4$ $49.3$ $66.3$ $0.86$ $0.79$ $1.00$ $0.47$ $0.70$ $0.84$ $0.61$ $0.75$ $0.24$
ADEPT-F $93.2$ $\underline{85.5}$ $69.9$ $56.3$ $76.2$ $0.50$ $0.50$ $0.75$ $0.53$ $0.80$ $0.68$ $0.65$ $0.63$ $0.13$
CCD-CDA $92.8$ 86.3 $65.3$ $50.7$ $73.8$ $0.79$ $0.79$ $0.83$ $0.80$ $0.70$ $0.79$ $0.83$ $0.79$ $0.29$
CCD-GPT 93.6 $85.1$ $70.4$ $\underline{56.3}$ $\underline{76.4}$ $0.50$ $0.43$ $0.58$ $0.60$ $0.35$ $0.53$ $0.61$ 0.51 0.07
CCD-Gemini $\underline{93.5}$ $83.6$ 72.9 56.3 76.6 $0.57$ $0.64$ $0.58$ $0.60$ $0.70$ $0.42$ $0.65$ $\underline{0.59}$ $\underline{0.11}$
We evaluate four models on all benchmarks, namely the original model (pre-trained with no explicit debiasing), the DPCE model, the ADEPT-F model, and our CCD.
Reducing Gender Biases In Table 1, our experiments reveal that our CCD with all data strategies (CDA, GPT, and Gemini) excels in debiasing, consistently outperforming baselines in the StereoSet and the CrowS-Pairs dataset and ranking second best in SEAT. It is worth noting that our method achieves a good ICAT score in the StereoSet dataset, indicating an excellent balance between performance and fairness. However, DPCE, while preserving great fairness, adversely affects its representation capability, as evidenced by a notably lower LMS score in the StereoSet dataset.
Preserving Representation Ability In Table 2, the GLUE results demonstrate that CCD-Gemini achieves the highest average performance, suggesting that it not only maintains but also enhances the model’s representational capabilities. Conversely, DPCE, which is intended to separate gender attributes from neutral text embeddings, adversely affects the model’s utility.
Bias in Information Retrival We conduct an evaluation using the Bias-IR dataset to assess the fairness of text embedding for long text sequences. Table 2 illustrates that our approach, CCD-Gemini and CCD-GPT outperforms others in fairness across numerous fields, as well as in average fairness.
Comparision of Data Strategy To demonstrate the effectiveness of our proposed data strategy, we conduct comparisons with CDA as shown in Table 1 and Table 2. Integrating our debiasing loss with both data strategies resulted in improved fairness. However, CDA experienced a drawback in maintaining model utility, due to potential grammatical inconsistencies introduced by its rule-based augmentation. This highlights the superiority of our LLM-based augmentation method. For different LLMs, we found that both ChatGPT and Gemini achieve good performance on fairness and utility.
Embedding Visualization Fig. 2 reveals that original BERT clusters male and female embeddings, indicating fairness issues. DPCE and ADEPT-F debias by separating sensitive attributes from neutral embeddings, harming the representation quality. Our CCD maintains the original embedding proximity and mixes male and female embeddings, achieving fairness effectively.
4 Conclusion
This paper presents a new method for generating fair text embeddings by ensuring conditional independence between sensitive attributes and text content. We show that conditional independence can be represented as content-conditioned equal distance in embedding spaces. By leveraging LLMs to augment datasets, we address the data scarcity issue and enable model debiasing. Extensive experiments in diverse datasets show that our approach enhances fairness while preserving utility.
Figure 2: T-SNE plots of embeddings that are processed by different methods. Our approach maintains embedding positions similar to BERT while mixing male and female embeddings thus achieving fairness.
5 Limitaions
In this study, we utilize gender bias to demonstrate the efficacy of our method. As our approach constitutes a general pipeline, we plan to extend our methodology to address other types of biases (e.g., race, age) in the future. Moreover, we discuss the application of our method in a binary gender setting, which generally does not reflect the real world where gender (and other biases) may not be strictly binary. Fortunately, our method is readily extensible to any number of dimensions. We consider this extension as part of our future work.
6 Ethical Consideration
Our work pioneers in mitigating biases in text embeddings, crucial for fairness and inclusivity in NLP applications. We introduce a method that ensures fair representation by achieving conditional independence between sensitive attributes and text embeddings, aiming to reduce societal biases. Employing LLMs for data augmentation represents ethical advancement in tackling inherent biases, moving towards equitable technology and inspiring future bias-aware research. Our contribution significantly advances AI fairness by validating a method that minimizes bias in text embeddings, promoting inclusivity in machine learning.
Appendix A Related Work
LLMs in creating dataset: Leveraging the success of large language models (LLMs), researchers have begun using them to generate various forms of training data, such as tabular data Borisov et al. (2022), relation triplets Chia et al. (2022), sentence pairs Schick and Schütze (2021), and instruction data Shao et al. (2023),et al.. As we focus on obtaining data with sensitive attribute information, data generation for text classification would be the most similar one among those applications. Recent efforts in generating data for text classification  Meng et al. (2022); Ye et al. (2022); Wang et al. (2019) primarily employ simple class-conditional prompts while focusing on mitigating issues of low quality after generation. However, these efforts encounter the challenge of inherent systematic biases present in LLMs Yu et al. (2023). Although recent study Yu et al. (2023) considered generated data bias, it focuses only on the diversity of topics and overlooks the inherent bias within words in a text (i.e., "child" occurs more frequently with "mother").
Debias Text Embedding: Bias in text embeddings (also known as sentence embedding) is a significant issue that arises when these models reflect or amplify societal stereotypes and prejudices found in their training data. To resolve the issue,  Liang et al. (2020) contextualizes predefined sets of bias attribute words to sentences and applies a hard-debias algorithm Bolukbasi et al. (2016). Contextualized bebiasing methods Kaneko and Bollegala (2021); Yang et al. (2023) apply token-level debiasing for all tokens in a sentence and can be applied at token- or sentence-levels Kaneko and Bollegala (2021) to debias pretrained contextualized embeddings. However, all the above methods aim to strictly achieve independence between text embedding and sensitive attributes, which may not achieve a good balance between fairness and utility.
Appendix B Algorithm Details
B.1 Notation
Table 3: Main notations used in this paper.
Basic Variables
$L$triangleq
loss function
$f$,$f^{ori}$triangleq
finetuned and original text embedding model.
$h$triangleq
Large language model.
$\theta_{p}$triangleq
Few-shot prompts that used to empower a LLM.
$A$, $a_{i}$triangleq
Sensitive attribute set and $i$-th sensitive attribute.
$S^{a_{i}}$,$S^{n}$triangleq
Text that relate to sensitive attribute $a_{i}$ and neutral text.
$C$,$C^{\prime}$triangleq
Content variable and predicted content.
$X^{a_{i}}$, $X^{n}$triangleq
words from group $a_{i}$ and neutral words in a text.
$V^{a_{i}}$triangleq
words list that contains all collected words related to attribute $a_{i}$.
B.2 Fainess Goals
Improving Fairness There are two goals in improving the fairness of a text embedding model Yang et al. (2023); Kaneko and Bollegala (2021). (1) Remove the superficial connection between the content $C$ and sensitive attributes $A$ so that the fairness is improved (2) Maintain the utility of the text embedding model so that the performance is protected.
Table 4: We utilize LLM to augment text into three gender categories: Male, Female, and Neutral. Below are sample examples of the generated data, where words containing gender information are highlighted in colors: red for male, blue for neutral, and orange for female.
Gender Generated Text
Male But because Rumsfeld wanted to prove a point about transforming strategy.
After championing the continuation of his hardline policy, his current strategy of negotiation is risky.
He has been very vocal in voicing discontent with the rule of Kirchner and that of his husband and predecessor, Néstor Kirchner.
Neutral But because the individual wanted to prove a point about transforming strategy.
After championing the continuation of their hardline policy, the current strategy of negotiation is risky.
They have been very vocal in voicing discontent with the rule of Kirchner and that of their spouse and predecessor, Néstor Kirchner.
Female But because Rachel wanted to prove a point about transforming strategy.
After championing the continuation of her hardline policy, her current strategy of negotiation is risky.
She has been very vocal in voicing discontent with the rule of Kirchner and that of her wife and predecessor, Néstor Kirchner.
B.3 Data Augmentation Details
We generated training data using the News-Commentary-v15 corpus Tiedemann (2012) focusing on gender bias. By employing Gemini and ChatGPT for data augmentation, we obtained datasets comprising 43,221 and 42,930 sample pairs, respectively. Each pair contains texts with identical content from male, female, and neutral perspectives.
B.4 Data Augmentation Results
In this section, we show data samples augmented by our algorithm, as described in Section 2.3, in Table 4. Words associated with sensitive information in the text are replaced with corresponding words from all different sensitive groups and their neutral counterparts. Ultimately, we obtained a dataset with texts of content $C$ from all groups $A$ and neutral.
B.5 Ommited Proofs
In this section, we give a detailed proof of definition 2.1.
Proof.
Firstly, we write the conditional independence for any $a_{i},a_{j}\in A$ :
$P(C^{\prime}|A=a_{i},C)=P(C^{\prime}|A=a_{j},C)=P(f(S^{n})|C)$ (6)
where $C^{\prime}$ is the content embedding. Then, assume an equal probability for different sensitive attributes $P(a_{1}|C)=...=P(a_{A}|C)$, which is true for a fair dataset that has balanced texts from all groups within content $C$ (can be obtained through our data augmentation strategy in Section 2.3). we can rewrite the Eq. 6 as:
$P(C^{\prime},A=a_{i}|C)=P(C^{\prime},A=a_{j}|C)$
$P(f(S^{a_{i}})|C)=P(f(S^{a_{j}})|C)$ (7)
Following Hinton and Roweis (2002); Yang et al. (2023) the probability that $C$ would pick $a_{i}$ as its neighbor (in other words, the probability $a_{i}$ conditioned on $C$):
$P_{f(S^{a_{i}})|C}=\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{i}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$ (8)
where $\rho$ is either set by hand or found by calculating the variance of $f(S^{a_{i}})-f(S^{n})$ over the whole training set (as in our experiments). Eq. 8 can be interpreted as follows: (1) Consider setting a Gaussian distribution with a covariance matrix equal to $\rho$ times the identity matrix at the embedding of a neutral text $S_{C}$ (with content $C$), which is denoted as $f(S_{C}^{n})$. Then, a text with the same content but containing sensitive information $a_{i}$ appears in the distribution with a probability proportional to $\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)$, represented as the numerator. (2) The denominator aggregates the aforementioned probabilities across all sensitive groups $a_{i}\in A$ and serves as the normalization factor. Then we combine the Eq. B.5 and Eq. 8 and obtain:
$\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{i}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$ $=\frac{\exp\left(-\frac{\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}{\sum_{a_{j}\in A}\exp\left(-\frac{\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)}$
$\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)$ $=\exp\left(-\frac{\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}}{2\rho^{2}}\right)$
$\lVert f(S_{C}^{a_{i}})-f(S_{C}^{n})\rVert^{2}$ $=\lVert f(S_{C}^{a_{j}})-f(S_{C}^{n})\rVert^{2}$ (9)
Thus we obtain the definition 2.1. As a result, achieving conditional independence between sensitive attributes and content embeddings is equivalent to achieving content-conditioned equal distance. ∎